OpenStudy (anonymous):

Hi, my problem is the following: "a engines factory separate from its daily production line of 350 pieces a sample of 30 items for inspection. The number of defective pieces is 14/day. What's the probability that the sample contains at least 3 defective pieces?" The exercise I've been doing points out that the correct answer is "0,108453", but what I find in any calculus I try is "0,9....". Can someone tell me where is my mistake?

OpenStudy (anonymous):

at least 3 means not none, not one, not two so computer those probabilities and subtract from one are you supposed to use binomial probability?

OpenStudy (anonymous):

\[\frac{14}{350}=\frac{1}{25}=.04\] if you are using binomial, probablity that none of the 30 are defective is \[.96^{30}\] probability one is defective is \[30\times .96^{29}.04\] and probability two are defective is \[\dbinom{30}{2}.96^{28}.4^2\] add these up and subtract from one

OpenStudy (anonymous):

I've used binomial probability, but the answer was 0,9....

OpenStudy (anonymous):

Tell me what you did find as result..

OpenStudy (anonymous):

let me check it

OpenStudy (anonymous):

not quite what the book has, but i like my answer just fine

OpenStudy (anonymous):

I've just found 0,1049, ignoring some digits. But it sounds plausible. :P Can you tell me why is it 14/350 instead of 14/30 if the sample is supposed to be 30?